OpenStudy (anonymous):

Find the formula for the Riemann sum obtained by dividing the interval [a,b] into n equal subintervals and using the right hand endpoint for each ck. then take a limit of these sums as n->approaches infinity to calculate the area under the curve over [a,b]. f(x) = 3x^2+2 over the interval [0,3]

OpenStudy (anonymous):

got an hour?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

the interval if broken up in to n parts, will each have length \(\frac{3}{n}\) so that is your \(\Delta x\)

OpenStudy (anonymous):

can we split the sum up in to two parts? that will make it tons easier

OpenStudy (anonymous):

well, I would prefer not. It is up to you.

OpenStudy (jhannybean):

\[f(x) = 3x^2 +2 \ , \ [0.3] \] for reimann's sums, you have to take the upper and lower limit of the function to get the most approximate area.

OpenStudy (anonymous):

lets to \(x^2\) and we can multiply the result by 3 at the end

OpenStudy (anonymous):

In general \(\Delta x = (b-a)/n\) and \(x_{iR} =a+\Delta x \times i \)

OpenStudy (anonymous):

\[x_0=0,x_1=\frac{3}{n},x_2-\frac{6}{n},x_3=\frac{9}{n}\] and in general \[x_k=\frac{3k}{n}\]

OpenStudy (anonymous):

then \[f(x_k)=(\frac{3k}{n})^2=\frac{9k^2}{n^2}\]

OpenStudy (anonymous):

When using right endpoint, \(x=0\) will not be used. You will have \(3/n \leq x \leq 3\)

OpenStudy (anonymous):

makes no difference

OpenStudy (jhannybean):

\[\ a = 0, \ b = 3\]\[\ \Delta x_i = x_i - x_{i-1} = \frac{b-a}{n}= \frac{3-0}{n} = \frac{2}{n}\]

OpenStudy (anonymous):

then \[\sum f(x_k)\Delta x=\sum (\frac{9k^2}{n^2})\frac{3}{n}=\frac{9}{n^3}\sum k^2\]

OpenStudy (anonymous):

ooops typo there \[\frac{27}{n^3}\sum k^2\]

OpenStudy (anonymous):

use the closed form of \[\sum_{k=1}^n k^2=\frac{n(n+1)(2n+1)}{6}\]

OpenStudy (anonymous):

let me know when I can show you what I got and tell me if it is correct

OpenStudy (jhannybean):

\[\ x_i = a+i \left( \frac{b-a}{n}\right) =\frac{3i}{n}\]\[\ Area_R=\sum_{i=1}^{n}f(x_i)(\Delta x_i) = \sum_{i=1}^{n}(3x^2+2)\left(\frac{3}{n}\right)=\sum_{i=1}^{n} \left[ 3\left(\frac{3i}{n}\right)^2 +2\right]\left(\frac{3}{n}\right)\]

OpenStudy (jhannybean):

Sure go ahead.

OpenStudy (jhannybean):

I'm getting \[\ \sum_{i=1}^{n} \left[ \frac{27i^2}{n^2} +2 \right]\left(\frac{3}{n}\right)\]\[\ \sum_{i=1}^{n} \left[ \frac{81i^2}{n^3}+ \frac{6}{n}\right]\]\[\ \frac{81}{n^3}\sum_{i=1}^{n} i^2 + \frac{6}{n}\sum_{i=1}^{n} 1\]

OpenStudy (jhannybean):

\[\frac{81}{n^3} \left[ \frac{n(n+1)(2n+1)}{6}\right] + \frac{6}{n}(n)\]

OpenStudy (jhannybean):

Whatever this reduces down to... one min

OpenStudy (jhannybean):

\[\ \frac{81(2n^3+3n^2+n)}{6n^3}\] this is what i'm getting.. but this is not fully reduced.

OpenStudy (anonymous):

ok tell me if this is right what I have here

OpenStudy (jhannybean):

Can you tell me how you got \(\ \frac{k}{n}\) as opposed to getting \(\ \frac{3k}{n}\)?

OpenStudy (anonymous):

what do you mean? i square k/n and then multiply it by 3

OpenStudy (jhannybean):

I meant to ask, where did the k/n come from in your equation? Can you tell me how you found it?

OpenStudy (jhannybean):

@satellite73 found the \(\ x_i\) the same method I had, which is \(\ \large\frac{b-a}{n}\). In your case, \(\ a= 0\) and \(b= 3\)

OpenStudy (anonymous):

I got it from the formula in the textbook

OpenStudy (anonymous):

\[(\frac{ 27i }{ n })+2*(3/n)\]

OpenStudy (jhannybean):

Yes, and the 3/n multiplies throughout. That is how I got the 81i/n^3. Understand? :o

OpenStudy (anonymous):

yes

OpenStudy (jhannybean):

Awesome \(^_^)/

OpenStudy (anonymous):

wait but I am confused about 6/n(n)

OpenStudy (jhannybean):

\[\ \sum_{i=1}^{n} \left[ \frac{27i^2}{n^2} +2 \right]\left(\frac{3}{n}\right)\]\[\ \sum_{i=1}^{n} \left[ \frac{81i^2}{n^3}+ \frac{6}{n}(1)\right]\]\[\ \frac{81}{n^3}\sum_{i=1}^{n} i^2 + \frac{6}{n}\sum_{i=1}^{n} 1\]\[\frac{81}{n^3} \left[ \frac{n(n+1)(2n+1)}{6}\right] + \frac{6}{n}(n)\]

OpenStudy (anonymous):

ok now what happens to the 6/n*n it reduces to 6? then what happens to the 6

OpenStudy (jhannybean):

Then you get \[\ \ \frac{81(2n^3+3n^2+n)}{6n^3}+6\]

OpenStudy (jhannybean):

You simplify the fraction then take the limit as n goes to \(\ \infty\)

OpenStudy (anonymous):

so when i take the limit don't I just substitute b?

OpenStudy (jhannybean):

I believe it's just \[\ \lim_{n \rightarrow \infty} \frac{81(2n^3+3n^2+n)}{6n^3}+6\] reduce the fraction, of course.

OpenStudy (anonymous):

\[\frac{ 54n^3+81n^2+27n }{ 2n^3 }+6\]

OpenStudy (anonymous):

840?

OpenStudy (jhannybean):

\[\ \lim_{n \rightarrow \infty} \frac{162n^3 + 243n+81n}{6n^3} + 6\]\[\lim_{n \rightarrow \infty} \left(\frac{162}{6} + \frac{243}{6n^2} +\frac{81}{6n^2} +6 \right)\] all the ones that have n's reduce to 0, and you add up all the remaining numbers.

OpenStudy (anonymous):

so it is not my simplification above?

OpenStudy (jhannybean):

your way is right as well, you just simplified \(\ \large \frac{81}{6} = \frac{27}{2}\) which is fine as well

OpenStudy (anonymous):

ok so does that simplify to\[30+81n^2+27n\]

OpenStudy (jhannybean):

So then you'd have \[\ \lim_{n \rightarrow \infty} \left(\frac{ 54n^3+81n^2+27n }{ 2n^3 }+6\right)\]\[\ \lim_{n \rightarrow \infty} \left( 27 + \frac{81}{n^2} + \frac{27}{2n^2} +6\right)\]and taking the limi of this you'd be left with 27+6 \[\ \ \lim_{n \rightarrow \infty} \left( 27 +\cancel{ \frac{81}{n^2}} + \cancel{\frac{27}{2n^2}} +6\right)\]\[\ = 27+6\]

OpenStudy (jhannybean):

Which is what I got in my previous method.

OpenStudy (jhannybean):

You need to understand if n is approaching infinity, the fractions with n in their denominator will approach 0.

OpenStudy (jhannybean):

As the numerator stays constant, the denominator gets exponentially large, therefore the fraction eventually becomes 0.

OpenStudy (anonymous):

arent I supposed to sub 3? and my simplification seems different.